OpenStudy (el_tucan):

Hello OpenStudy Friends! WITHOUT using a calculator, find one critical value for the function y = tan(x)

OpenStudy (el_tucan):

I know I have to find the first derivative, which is y' = 1 + sec^2(x).

OpenStudy (el_tucan):

I then try setting it to zero and move the 1 over making it negative.

OpenStudy (el_tucan):

from there I'm lost...

OpenStudy (anonymous):

y' of tan(x) is (sec^2(x)) not (1+ sec^2(x))

OpenStudy (el_tucan):

oh sorry, it's y = x + tan(x)

jimthompson5910 (jim_thompson5910):

sec^2(x) = -1 sec(x) = sqrt(-1) so this implies that there are no real numbers x that make sec^2(x) + 1 = 0 true

OpenStudy (el_tucan):

looks good to me.

OpenStudy (el_tucan):

would finding the inverse value help anything? what does that do anyway?

jimthompson5910 (jim_thompson5910):

well the inverse secant of sqrt(-1) isn't defined (I don't think anyways) it could be defined in some strange way, but idk

jimthompson5910 (jim_thompson5910):

if you're only considering real numbers for x, then there are no solutions to sec^2(x) + 1 = 0 so there are no critical numbers

OpenStudy (el_tucan):

awesome thnx

jimthompson5910 (jim_thompson5910):

np